11) The standard error of the sample mean is equal to 5 when n=25. If the sample size increases by a factor of four, who will the standard error change?
A.

It will double.
B.

It will be cut to ¼ of 5.
C.

It will quadruple.
D.

It will be cut in half.

Answer by stanbon(75887)

11) The standard error of the sample mean is equal to 5 when n=25. If the sample size increases by a factor of four, how will the standard error change?
---
E = t*s/sqrt(n)
---
If sample size becomes 4n you get:
E = t*s/sqrt(4n) = ts/(2sqrt(n))
---
Cross-multiply and you get
2E = ts/sqrt(n)
-------
Answer: A
==========================
Cheers,
Stan H.
==========================

A.
It will double.
B.
It will be cut to ¼ of 5.
C.
It will quadruple.
D.
It will be cut in half.